# Are Small Language Models Ready to Compete with Large Language Models for Practical Applications?

### v2

@@ 1 Introduction
The field of NLP has advanced significantly with the rise of Language Models (LMs). Rapid development of LMs Brown et al. (2020); Touvron et al. (2023); Almazrouei et al. (2023); Team et al. (2024b); DeepSeek-AI (2024) has expanded their use across numerous types like Title Generation (Keles and Bayraklı, 2024), Data Exploration (Ma et al., 2023), Dialogue act recognition (Qiang et al., 2024); domains like Economics & Finance (Rajpoot et al., 2024; Yu et al., 2023), Politics (Feng et al., 2023), Nutrition & Food (Yang et al., 2024), News (Kuila and Sarkar, 2024); and reasoning types (Huang and Chang, 2023) like ANALOGICAL (Wijesiriwardene et al., 2023) and Multi-hop (Pan et al., 2021) reasoning.
@@
Initially, LMs were relatively weak like GPT-2 (Radford et al., 2019), too large in size like GPT-3 (Brown et al., 2020), expensive like GPT-4 (OpenAI et al., 2024), and/or closed and accessible only via APIs. These pose challenges of accessibility, computational resource limitations (Ding et al., 2024), costs (Jimenez Gutierrez et al., 2022) (GPT-4 Turbo charges \$10/1M input, \$30/1M output tokens), data privacy, inflexibility of customization, and so on. However, there has been recent rise in competitive LMs which are relatively small and openly available. These properties remove the aforementioned limitations, and provide additional benefits like on-device usage, faster inference time, easier compliance and security management.
@@
But do these smaller LMs fit every use case like the SOTA LLMs? They vary a lot in terms of training data, pre-training strategies, and architectural decisions. Further, utilization strategies like zero-shot usage, customizing pre-trained models (e.g., fine-tuning (Mosbach et al., 2023)), using in-context learning (Wei et al., 2022a; Dong et al., 2023), prompt engineering (Brown et al., 2020) can differ with varying need. Writing helpful prompts also requires time and domain expertise. For someone to use these small, open LMs, they need to conduct their analysis with constraints of time, money, computational resources is a complicated task, and identify the LMs fit for their use. Technical reports of some LMs (Team et al., 2024b, c) report performance on different benchmarks, but they are more theoretical than in a practical usage setting. Therefore, there is a need for a comprehensive evaluation of these LMs and determining their capabilities for different use-cases in different practical applications and determine if one can use them instead of SOTA LLMs within tolerable levels of performance degradation.
To bridge this gap, we perform this extensive, in-depth experimental analysis with 10 openly available LMs between 1.7B–11B parameters. We also vary prompt styles by using chat-style task definitions, in-context examples that are closer to how users interact with LMs, and evaluate them across three aspects – task types, application domains and reasoning types, which can be considered as three different ways of categorizing an application area aiming to use LMs. We propose a schema by selecting 12, 12, and 10 entities from each aspect respectively in English language covering a broad range of areas, and group similar entities. For example, in application domains, we group ‘Social Media’ and ‘News’ in ‘Media and Entertainment’. This three-tier structure (aspect, group, entity) allows finding patterns in capabilities of LMs at multiple level, along different aspects.
A recent work (Zhao et al., 2024) dives into efficacy of LoRA (Hu et al., 2022a) fine-tuning of smaller LMs, but uses static prompts, and limits to varying task types only. Further, usage strategy of LMs can significantly vary – fine-tuning (Mosbach et al., 2023), PEFT (Han et al., 2024), multi-stage pipeline based, or directly with/without prompt engineering due to lack of sufficient data, resources, or need for customization. Hence, we consider semantic correctness of outputs as a measure of LMs’ innate ability, and evaluate 5 pre-trained and 5 instruction-tuned (IT) (Ouyang et al., 2022) LMs out-of-the-box with 8 prompt styles. We show that scale can be countered with careful selection of small, open LMs, which can enable matching and outperforming models like GPT-4o-mini, GPT-4o (OpenAI, 2023), DeepSeek-v2 (DeepSeek-AI, 2024) and Gemini-1.5-Pro (Team et al., 2024a) for practical applications. Our proposed framework enables this analysis and identifies patterns in strengths and weaknesses at 3 hierarchical levels.
@@
(ii) Conduct an in-depth experimental analysis of semantic correctness of outputs of 10 open, small LMs in 2B–11B size based on the framework.
(iii) Show that appropriate selection of open, small LMs can lead to outperforming SOTA LLMs like GPT-4o-mini, Gemini-1.5-Pro and DeepSeek-v2, and competing with GPT-4o.
@@ 2.4 Evaluation Metrics
Evaluating LM outputs using n-gram metrics like ROUGE (Lin, 2004), METEOR (Banerjee and Lavie, 2005), etc., have limitations in terms of coherence, consistency, relevance, fluency of those metrics as discussed in recent works (Sai et al., 2021; Xiao et al., 2023) in depth. To address this, we evaluate LM’s knowledge via semantic correctness of outputs using BERTScore (Zhang et al., 2019) recall with roberta-large (Liu et al., 2019) which greatly limits these issues.
Some tasks, like classification, aren’t generation tasks, but we still consider them as one since gives a uniform evaluation paradigm. By aligning outputs using fine-tuning/ICL (Zhao et al., 2023), verbalizers (Hu et al., 2022b), post-processing, labels can be obtained from language outputs.
@@ 2.5 Language Models Used
To cover a broad range of SOTA small, open LMs across sizes, families, our experiments utilize Gemma-2B, Gemma-7B (Team et al., 2024b), Llama-3-8B Touvron et al. (2023); AI@Meta (2024), Mistral-7B-v0.3 Jiang et al. (2023), and Falcon-2-11B Almazrouei et al. (2023); TIIUAE (2024). We also take their instruction-tuned (IT) versions (except Falcon-2-11B - not available). But, we omit Mistral-7B pre-trained from main paper’s discussion as its results weren’t competitive, and Gemma-2 series (Team et al., 2024c) since their performance was below Gemma. Model and implementaton details are discussed more in Appendix C,  G. In this paper, suffix "-I" indicates instruction-tuned.
@@ 3.2 Comparison Across Task Types
In IT models, Mistral-7B-I performs best on all task types, with Gemma-2B-I and SmolLM-1.7B-I competing for the second-best. At group level, we find the difference to be smaller for linguistic relationship and generation tasks, but large for semantic & pragmatic analysis tasks. Like their pre-trained variants, Gemma-7B-I and Llama-3-8B-I seldom compete with Gemma-2B-I in some tasks, but never outperforms it. So, Gemma-2B, SmolLM-1.7B-I and Mistral-7B-I can be selected based on performance and resources trade-offs.
@@ 3.3 Comparison Across Application Domains
Particularly for pre-trained models, the performance is very sensitive across domains. For social sciences & humanities, and science & technology domain groups, Falcon-2-11B performs the best with Gemma-2B and Llama-3-8B following. Gemma-2B and Falcon-2-11B are not always the best ones. In health and medical tasks, Gemma-7B outperforms all models. Falcon-2-11B and Gemma-2B suffer a significant performance degradation in this group. Therefore, for domains, the choice of pre-trained LMs depends on the use case and other constraints. SmolLM-1.7B felt like a strong choice in task types, but here we see here that it struggles with these domains. It’s strength in Section 3.2 might be from other domains not considered here, showing its sensitivity with domains.
@@ 3.5 Comparison with State-of-the-art LLMs
We compare small, open LMs Gemma-2B, Falcon-11B, Mistral-7B-I and Gemma-2B-I (overall best two from each category) with recent SOTA LLMs like GPT-4o-mini, GPT-4o (OpenAI et al., 2024; OpenAI, 2023), DeepSeek-v2 (DS-2) (DeepSeek-AI, 2024) and Gemini-1.5-Pro (Team et al., 2024a). GPT-4o, Gemini-1.5-Pro and GPT-4o-mini are costly, large, closed models accessible using APIs. DeepSeek-v2 is open, but has 243B parameters. We use 8 examples with task definition for SOTA models, and report results in Figure 6.
@@ 4 High-level Takeaways
For the LMs we experimented with, among pre-trained models, we recommend using Gemma-2B and Falcon-2-11B based on different aspects and entities, but sometimes, Gemma-7B, Llama-3-8B can be great for some application domains (also detailed in Appendix B). The performance of pre-trained models can be taken as a measure of their knowledge of different use-cases. Based on other factors like availability, compliance, size, right LM can be selected and customized as needed. Limitations of some pre-trained models are discussed in Appendix F.2.
@@
We also study the performance trade-off for multiple prompt styles and recommend the best one for these models. As noted in Section 3.6, having a chat-style task description to guide the LM is recommended. But, having more in-context examples is not always better, and considering use-case and LMs, the right number of example can vary. The models are also robust to changes in task definitions, if it can provide all (or most) information to complete the task. They are also reasonably robust to subtle intended/unintended incorrectness in definitions, which is analyzed in Appendix D.4.
@@ 5 Conclusion
We identify some limitations of using SOTA, proprietary LLMs and show that open LMs with 1.7B–11B parameters can be effective for applications. We create a three-tier evaluation framework and analyze semantic correctness of output of 10 LMs across multiple hierarchical umbrellas. Using this framework, we demonstrate that while these models don’t work best in every scenario, if selected properly, they are effective and can compete with and outperform models like Gemini-1.5-Pro, DeepSeek-v2 and GPT-4o-mini and GPT-4o despite being 10-20 folds smaller in size. We also provide a guide in Appendix A on how one can this work to select an LM for one’s specific needs. We hope that our contributions will enable the community to make a confident shift towards considering using these small, open LMs for their need.

### v3

@@ 1 Introduction
The field of NLP has advanced significantly with the rapid development of Language Models (LMs) Brown et al. (2020); Touvron et al. (2023); Almazrouei et al. (2023); Team et al. (2024b); DeepSeek-AI (2024), which has expanded their use across numerous types like Title Generation (Keles and Bayraklı, 2024), Data Exploration (Ma et al., 2023), Dialogue act recognition (Qiang et al., 2024); domains like Economics & Finance (Rajpoot et al., 2024; Yu et al., 2023), Politics (Feng et al., 2023), Nutrition & Food (Yang et al., 2024), News (Kuila and Sarkar, 2024); and reasoning types (Huang and Chang, 2023) like ANALOGICAL (Wijesiriwardene et al., 2023) and Multi-hop (Pan et al., 2021) reasoning.
@@
Despite the growing variety of LMs, their usage in downstream applications is heavily skewed towards limited ones. Analyzing around 50 papers from 2024, we found that while 82.3% of methods utilized GPT-family LMs, only 41.1% used Llama variants, and less than 11.8% experimented with other alternatives like Mistral (Jiang et al., 2023) and Falcon (Almazrouei et al., 2023). Some studies also report issues like garbage output and hallucinations (Alhamed et al., 2024), but domain experts often lack the tools to address them effectively through informed LM choice, or correct ways to prompt them. Even the methods that experiment with multiple LMs often select models without a strong motivation (Kuila and Sarkar, 2024).
Apart from performance, many of the new LMs are smaller in size, and openly available. Despite the undeniable success of large, proprietary LMs like GPT-4 (OpenAI et al., 2024) and Llama-2 70B (Touvron et al., 2023), their inaccessibility due to limited API access, high costs (Jimenez Gutierrez et al., 2022), concerns around data privacy (for GPT), and massive computational demands (Ding et al., 2024) (for Llama) pose significant barriers of usage. Small, open LMs can navigate around those, and also provide additional benefits like on-device usage, faster inference time, data privacy, easier compliance and security management, and low-cost maintainance. For many practitioners -— especially those in research, startups, or sectors with limited resources or high security risk – leveraging these presents an appealing alternative for functional, financial, or business reasons.
@@
But these new, small, LMs vary a lot in terms of training data, pre-training strategies, and architectural decisions. Additionally, they may not perform globally well like SOTA LLMs due to limitations of scale (Kaplan et al., 2020). Utilization strategies of LMs in inference pipelines can also differ, like zero-shot usage, customizing pre-trained models (e.g., fine-tuning (Mosbach et al., 2023)), using in-context learning (Wei et al., 2022a; Dong et al., 2023), prompt engineering (Brown et al., 2020). Writing effective prompts also requires time and domain expertise. So, users need to conduct thorough analysis before choosing the right LM and usage strategy within constraints of time, money, computational resources, which is a complicated task. Although technical reports of some LMs (Team et al., 2024b, c) provide some insights, not all of them capture real-world, practical scenarios. Therefore, there is a need for a comprehensive practical evaluation framework which can enable determining capabilities of LMs in multiple practical applications, and effective ways to prompt them.
To bridge this gap, we propose a comprehensive framework for evaluating LMs in practical settings along three aspects: task types, application domains, and reasoning types. For each aspect, we select 12, 12, and 10 entities in English, grouping similar ones (e.g., ’Social Media’ and ’News’ under ’Media and Entertainment’). This three-tier structure (aspect, group, entity) helps identifying patterns in LM capabilities across multiple levels. Using Super-Natural Instructions (Wang et al., 2022), a meta-dataset encompassing various NLP benchmarks, we evaluate LMs on task instances within this framework. LM usage strategies vary significantly – ranging from fine-tuning (Mosbach et al., 2023), PEFT (Han et al., 2024) or direct usage with/without prompt engineering. Thus, we assess semantic correctness of outputs as an indicator of LMs’ inherent abilities, evaluating five pre-trained and five instruction-tuned (IT) (Ouyang et al., 2022) models across eight prompt styles. Our results show that with careful selection, impact of scale can be reduced. Correctly chosen small, open LM can rival and even outperform models like GPT-4o-mini, GPT-4o (OpenAI, 2023), DeepSeek-v2 (DeepSeek-AI, 2024), and Gemini-1.5-Pro (Team et al., 2024a), while providing additional benefits. We also evaluate LMs with paraphrases of task definitions to show that results are robust against dataset-induced biases.
@@
(ii) Conduct an in-depth experimental analysis of semantic correctness of outputs of 10 open, small LMs in 1.7B–11B size based on the framework.
(iii) Show that appropriate selection of open, small LMs can lead to outperforming SOTA LLMs like GPT-4o-mini, Gemini-1.5-Pro, and competing with GPT-4o.
@@ 2.4 Evaluation Metrics
As per the analysis of recent works (Sai et al., 2021; Xiao et al., 2023), evaluating LM outputs using n-gram metrics like ROUGE (Lin, 2004), METEOR (Banerjee and Lavie, 2005), etc., have limitations in terms of coherence, consistency, relevance, and fluency. These works also show that BERTScore-recall (Zhang et al., 2019) limits this to a great extent. To be consistent, we evaluate LM’s knowledge via semantic correctness of outputs using BERTScore (Zhang et al., 2019) recall with roberta-large (Liu et al., 2019).
Some tasks, like classification, aren’t generation tasks, but we still consider them as one since they give a uniform evaluation paradigm. By aligning outputs using fine-tuning/ICL (Zhao et al., 2023), verbalizers (Hu et al., 2022), post-processing, labels can be obtained from language outputs.
@@ 2.5 Language Models Used
To cover a broad range of SOTA small, open LMs across sizes, families, our experiments utilize Gemma-2B, Gemma-7B (Team et al., 2024b), Llama-3-8B Touvron et al. (2023); AI@Meta (2024), Mistral-7B-v0.3 Jiang et al. (2023), and Falcon-2-11B Almazrouei et al. (2023); TIIUAE (2024). We also take their instruction-tuned (IT) versions (except Falcon-2-11B - not available). But, we omit Mistral-7B pre-trained from discussion as its results weren’t competitive, and Gemma-2 series (Team et al., 2024c) since their performance was below Gemma. Model and implementation details are discussed more in Appendix C,  G. In this paper, suffix "-I" indicates instruction-tuned.
@@ 3.2 Comparison Across Task Types
In IT models, Mistral-7B-I performs best on all task types, with Gemma-2B-I and SmolLM-1.7B-I competing for the second-best. At group level, we find the difference to be smaller for linguistic relationship and generation tasks, but large for semantic & pragmatic analysis tasks. Like their pre-trained variants, Gemma-7B-I and Llama-3-8B-I seldom compete with Gemma-2B-I in some tasks, but never outperform it. So, Gemma-2B, SmolLM-1.7B-I and Mistral-7B-I can be selected based on performance and resources trade-offs.
@@ 3.3 Comparison Across Application Domains
Particularly for pre-trained models, the performance is very sensitive across domains. For social sciences & humanities, and science & technology domain groups, Falcon-2-11B performs the best with Gemma-2B and Llama-3-8B following. Gemma-2B and Falcon-2-11B are not always the best ones. In health and medical tasks, Gemma-7B outperforms all models. Falcon-2-11B and Gemma-2B suffer a significant performance degradation in this group. Therefore, for domains, the choice of pre-trained LMs depends on the use case and other constraints. SmolLM-1.7B felt like a strong choice in task types, but here we see that it struggles with these domains. Its strength in Section 3.2 might be from other domains not considered here, showing its sensitivity with domains.
@@ 3.5 Comparison with State-of-the-art LLMs
We compare small, open LMs Gemma-2B, Falcon-11B, Mistral-7B-I and Gemma-2B-I (overall best two from each category) with recent SOTA LLMs like GPT-4o-mini, GPT-4o (OpenAI et al., 2024; OpenAI, 2023), and Gemini-1.5-Pro (Team et al., 2024a). GPT-4o, Gemini-1.5-Pro and GPT-4o-mini are costly, large, closed models accessible using APIs. We use 8 examples with task definition for SOTA models, and report results in Figure 6.
@@ 4 High-level Takeaways
For the LMs we experimented with, among pre-trained models, we recommend using Gemma-2B and Falcon-2-11B based on different aspects and entities, but sometimes, Gemma-7B, Llama-3-8B can be great choices (also detailed in Appendix B). The performance of pre-trained models can be taken as a measure of their knowledge of different use-cases. Based on other factors like availability, compliance, size, right LM can be selected and customized as needed. Limitations of some pre-trained models are discussed in Appendix F.2.
@@
We also study the performance trade-off for multiple prompt styles and recommend the best one for these models. As noted in Section 3.6, having a chat-style task description to guide the LM is recommended. But, having more in-context examples is not always better, and considering use-case and LMs, the right number of example can vary. The models are also robust to changes in task definitions, if it can provide all (or most) information to complete the task. They are also reasonably robust to subtle intended/unintended incorrectness in definitions, which is analyzed in Appendix D.4. In appendix B, we also compare and show that the small LMs even outperform DeepSeek-v2 (DS-2) (DeepSeek-AI, 2024) in many categories.
@@ 5 Conclusion
We identify some limitations of using SOTA, proprietary LLMs and show that open LMs with 1.7B–11B parameters can be effective for applications. We create a three-tier evaluation framework and analyze semantic correctness of output of 10 LMs across multiple hierarchical umbrellas. Using this framework, we demonstrate that while these models don’t work best in every scenario, if selected properly, they are effective and can compete with and outperform models like Gemini-1.5-Pro, GPT-4o-mini and GPT-4o despite being 10-20 folds smaller in size. We also provide a guide in Appendix A on how one can this work to select an LM for one’s specific needs. We hope that our contributions will enable the community to make a confident shift towards considering using these small, open LMs for their need.